OpenStudy (cailindevereux):

Use an Addition or Subtraction Formula to write the expression as a trigonometric function of one number. cos 10° cos 80° − sin 10° sin 80° also the exact answer

zepdrix (zepdrix):

\[\large\rm \cos\color{indianred}{A}\cos\color{royalblue}{B}-\sin\color{indianred}{A}\sin\color{royalblue}{B}\quad=\quad \cos(\color{indianred}{A}+\color{royalblue}{B})\]

zepdrix (zepdrix):

\[\large\rm \cos\color{indianred}{10}\cos\color{royalblue}{80}-\sin\color{indianred}{10}\sin\color{royalblue}{80}\quad=\quad ?\]

Directrix (directrix):

@cailindevereux Look at the identity posted by @zepdrix Answer this: If A = 10° and B = 80°, then A + B = what?

Directrix (directrix):

After you get the sum of A and B, what is the value of the cosine of that number of degrees?